### v1

# Some Structures arising from the Farey Fractal

Abstract
When one study the shape of Farey sequences, that is the shape of the binary planar rooted tree
$C(R)=\{(x,y)\in\mathbb{N}\times\mathbb{N}\;\mid\;\gcd(x,y)=1,\;x+y\leq R\},$
one rediscovers the Zeckendorf expansion of $R$ as a sum of non-consecutive Fibonacci numbers.
With $\varphi$ the golden ratio, we notice the subtle difference between identifying
$\varphi^{2}\;\longleftrightarrow\;\varphi+1,\quad\text{hence}\quad\varphi^{n+1}\;\longleftrightarrow\;\varphi^{n}+\varphi^{n-1},$
obtaining the totally disconnected, locally compact (pro finite) ”Fibonadic numbers”,
versus identifying also
$\varphi\;\longleftrightarrow\;\sum_{j\geq 0}\varphi^{-2j},\quad\text{hence}\quad\varphi^{n+1}\;\longleftrightarrow\;\sum_{j\geq 0}\varphi^{\,n-2j},$
obtaining the continuum of non-negative real numbers.
This work explores new arithmetic and combinatorial structures arising from the interplay between Farey-type graphs, Fibonacci expansions, and operadic constructions. We introduce Fibonadic numbers, defined as an inverse limit under the Zeckendorf shift, equipped with a metric, order, and commutative rig structure. A normalization lemma provides canonical representatives, while quotients of the non-zero Fibonadic numbers under shifts and it’s fundamental domain covering the circle via $\varphi$–values. Levels with associated functions $f^{z}$ encode the decomposition of $\mathcal{X}$ into arithmetic layers. This research links number theory, combinatorics, discrete dynamics, operads, fractal geometry and the golden ratio.
Put
$\mathbb{N}^{+}=\{1,2,3,\ldots\},\qquad\mathbb{N}^{++}=\{2,3,4,\ldots\}.$
Idea: instead of studying
$\mathbb{N}^{++}\;\backslash\;\mathbb{N}^{++}\cdot\mathbb{N}^{++}\;=\;\text{Primes},$
study
[…]
That is,
$\mathcal{X}=\{\,(x,y)\in\mathbb{N}^{+}\times\mathbb{N}^{+}\;\mid\;\gcd(x,y)=1\,\}\equiv\;\mathbb{Q}^{+}$
$(x,y)\;\longleftrightarrow\;\frac{y}{x}$
Thus $\mathcal{X}$ is the free abelian group on the set of primes, but it has more structure!
Fibonacci numbers
$a_{0}=0,\;a_{1}=1,\;a_{2}=1,\;a_{3}=2,\;a_{4}=3,\;a_{5}=5$
$a_{n+1}=a_{n}+a_{n-1},\;all\;n$
Binet formula
$a_{n}=\frac{\varphi^{n}-\psi^{n}}{\sqrt{5}}$
with $\varphi$ the golden ratio:
$\varphi=\frac{1+\sqrt{5}}{2},\quad\varphi^{2}=\varphi+1$
and with $\psi$ its conjugate
$\quad\psi=\frac{1-\sqrt{5}}{2}\quad=\quad 1-\varphi=-\frac{1}{\varphi}$
Put $f_{-n}$ = 0 for all $n\geq 0$
$f_{0}=0,\;f_{1}=1,\;f_{2}=2,\;f_{3}=3,\;f_{4}=5,\;f_{5}=8,\;f_{6}=13$
$f_{n+1}=f_{n}+f_{n-1},\quad n>1$
$f_{n}=a_{n+1},\quad n>0$
Zeckendorf expansion
Let $R\in\mathbb{N}^{+}$.
$M(R)=\max\{n\mid f_{n}\leq R\}$
$\delta R=R-f_{M(R)}$
$\Rightarrow\quad R=\sum_{j\geq 0}f_{M(\delta^{\,j}R)}$
Define
$Z(R)\in\{0,1\}^{\mathbb{N}^{+}},\quad Z(R)_{n}=1\iff n=M(\delta^{\,j}R)\text{ for some }j.$
$R=\sum_{n\geq 1}Z(R)_{n}\,f_{n}=\sum_{n\geq 1}z_{n}f_{n}.$
Note: The sequence $z_{n}$ has no consecutive 1’s, $z_{n}\bullet z_{n+1}\equiv 0$ for all n
Definitions
$\pi(R)=\sum_{n\geq 1}z_{n}f_{n-1},\qquad j(R)=\sum_{n\geq 1}z_{n}f_{n+1}$
$\pi\circ j=id$
$Z(\pi(R))_{n}=Z(R)_{n+1}$
$Z(j(R))_{n}=\begin{cases}0&n=0,\\ Z(R)_{n-1}&n>0\end{cases}$
Thus $\pi$ erases the last digit of the Zeckendorf expansion of a number, and $j$ adds a zero to the expansion.
Note that the maps $\pi$ and $j$ do not preserve addition, but their coboundaries take values in $\{-1,0,1\}$.
Lemma
$\partial\pi(x,y)=\pi(x+y)-\pi(x)-\pi(y)\in\{-1,0,1\}$
$\partial j(x,y)=j(x+y)-j(x)-j(y)\in\{-1,0,1\}$
Stern–Brocot Tree and Farey Graph
Let
$SL_{2}(\mathbb{N})=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\,\middle|\,ad-bc=1,\;a,b,c,d\in\mathbb{N}\right\}.$
It is the free monoid on two generators
$\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\quad\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$.
We have a bijection which we use as identification.
$SL_{2}(\mathbb{N})\quad\longleftrightarrow\quad X=\left\{(x,y)\in\mathbb{N}^{2}\;|\;\gcd(x,y)=1\right\}$
$g\longmapsto(1,1)g,\quad g_{v}=\begin{pmatrix}v^{-}\\ v^{+}\end{pmatrix}\;\longmapsfrom\;v$
[…]
The planar binary tree structure, with vertices $X$ and root $(1,1)$, is obtained by viewing $v+v^{+}>v+v^{-}$ as the descendants of $v$, and $v$ is the Mother of $v+v^{+}$ and $v+v^{-}$.
Farey Graph
The oriented graph’s structure is given by viewing each $g\in X$ as an edge going from the lower bound $(1,0)g$ to the upper bound $(0,1)g$. It has edges $X$, and vertices
$\overline{X}\;=\;X\,\sqcup\,\{\,0:=(1,0),\;\infty:=(0,1)\,\}.$
Note that for $v\neq(1,1)$, among $v^{+},v^{-}$ there is a longer vector, the Mother of $v$ denoted $M(v)$, and a smaller one, called the Father of $v$ and denoted $F(v)$.
Finite subtrees
Let $\mathcal{C}=$ finite subtrees of $X\quad\Longleftrightarrow\quad$ paths in $\bar{X}$ from 0 to $\infty$.
$c\;\;\Longleftrightarrow\;\;\partial c=\{\,0=(1,0)\,\}\;\sqcup\;c\;\sqcup\;\{\,\infty=(0,1)\,\}.$
$c\in\mathcal{C}[n]$ if #c=n1\#c=n-1, $\partial c=n+1$, and $\partial c$ has $n$ steps.
$\mathcal{C}=\bigcup_{n\geq 1}\mathcal{C}[n]$
For $c\in\mathcal{C}$, let $\Phi c$ denote the set of branching points of $c$:
$\Phi c\;=\;\{\,v\in c\;\mid\;v+v^{+}\in c\ \text{and}\ v+v^{-}\in c\,\}.$
We concentrate on
$C(R)=\{(x,y)\in X\;\mid\;|(x,y)|:=x+y\leq R\}$
and on
$C(R,n)\;:=\;\Phi^{n}C(R)=\Bigl{\{}v\in X\;\Big{|}\;|v|_{n}:=a_{n+2}|M(v)|+a_{n+1}|F(v)|\leq R\Bigr{\}}.$
Note that the paths in the tree X outgoing from the root, that go fastest out are $\{(f_{n},f_{n+1})\}$ and $\{(f_{n+1},f_{n})\}$
Thus:
$C(R,M(R)-1)=\varnothing,\qquad C(R,M(R)-2)=\{(1,1)\}.$
There is an operad structure on $\mathcal{C}$, and
$C(R,n)=C(R,n+1)\cdot\lambda^{R}_{n}$
with “d.n.a”
$\lambda^{R}_{n}=\sum_{j}\lambda^{R}_{n,j}\;\bullet\;\bigl{[}c_{j},c_{j+1}\bigr{]}\in\mathbb{Z}\;\partial C(R,n+1).$
That is, for $\bigl{[}c_{j},c_{j+1}\bigr{]}\in\partial C(R,n+1)$ and $\lambda_{j}=\lambda^{R}_{n,j}\in\mathbb{Z}$
to construct $\partial C(R,n)$ from $\partial C(R,n+1)$ one replaces the step
$c_{j},c_{j+1}]$
by
$\begin{cases}c_{j}\;\longrightarrow\;[c_{j}+c_{j+1}]\;\longrightarrow\cdots\longrightarrow[c_{j}+(\lambda_{j}+1)c_{j+1}]\longrightarrow c_{j+1},&\lambda_{j}>0,\\[6.0pt] c_{j}\;\longrightarrow[c_{j}+c_{j+1}]\longrightarrow c_{j+1},&\lambda_{j}=0\quad\text{(just add mediant)},\\[6.0pt] c_{j}\;\longrightarrow[c_{j+1}+(|\lambda_{j}|+1)c_{j}]\;\longrightarrow\cdots\longrightarrow[c_{j+1}+c_{j}]\longrightarrow c_{j+1},&\lambda_{j}<0.\end{cases}$
The $n$-th layer
Define the $n$-th layer:
$\ell(n,R):=C(R,M(R)-n-1)\setminus C(R,M(R)-n)$
$\{v\in X\;\mid\;|v|_{M(R)-n-1}\leq R<|v|_{M(R)-n}\}.$
and the base
$b(n,R)=C(R,M(R)-n-1)=\bigsqcup_{1\leq m\leq n}\ell(m,R)$
$\left\{\,v\in X\;\middle|\;f_{M(R)-n}\;\big{|}M(v)\big{|}+f_{M(R)-n-1}\;|F(v)|\leq R\,\right\}.$
Note that for fixed $v$ the norms $|v|_{k}$ satisfy the same recurrence
$|v|_{k+1}=|v|_{k}+|v|_{k-1},$
hence
$\pi(|v|_{k+1})=|v|_{k},$
for $k$ large $\bigl{(}f_{k}>|v|^{2}\bigr{)}$, and since $\pi$ is order-preserving,
[…]
Taking k=M(R)nk=M(R)-n, we get for $M(R)\gg n$,
$\ell(n,R)=\ell(n,\pi(R)).$
Fibonadic Numbers
Define the (non–negative) Fibonadic Numbers $\mathcal{F}$ as the inverse limit of $\mathbb{N}$ with respect to $\pi$:
$\mathcal{F}:=\lim_{\longleftarrow}\bigl{\{}\mathbb{N}\xleftarrow{\;\pi\;}\mathbb{N}\xleftarrow{\;\pi\;}\mathbb{N}\xleftarrow{\;\pi\;}\cdots\bigr{\}}.$
Alternatively, using Zeckendorf expansion:
$\mathcal{F}=\Bigl{\{}z=(z_{n})\in\{0,1\}^{\mathbb{Z}}\;\Big{|}\;z_{n}=0\text{ for }n\gg 0,\;z_{n}z_{n-1}=0\text{ for all }n\Bigr{\}}.$
For $z\neq z^{\prime}$, let
$\delta(z,z^{\prime}):=\max\{\,n\;\mid\;z_{n}\neq z^{\prime}_{n}\,\}.$
Then $d(z,z^{\prime}):=\varphi^{\delta(z,z^{\prime})}$ is a metric, and $\mathcal{F}$ is complete.
Put $z<z^{\prime}$ if for $n=\delta(z,z^{\prime})$, $z_{n}=0$, $z^{\prime}_{n}=1$. It’s a total order on $\mathcal{F}$.
For $z=(z_{n})\in\mathcal{F}$, let $\{R^{z}_{n}\}_{n\in\mathbb{N}}$ be the associated $\pi$–coherent sequence:
$R^{z}_{n}:=\sum_{j}z_{j}\,f_{n+j}\;\;\in\mathbb{N},$
and
$\pi(R^{z}_{n+1})=R^{z}_{n}.$
We get a decomposition of $\mathcal{X}$ into levels:
$\mathcal{X}=\bigsqcup_{n\geq 1}\ell(n,z),$
with
$\ell(n,z)\equiv\ell(n,R^{z}_{m})\quad\text{for any $m\gg n$}.$
and base
$b(n,z)=\bigsqcup_{1\leq m\leq n}\ell(m,z)$
We let the levels function
$f^{z}:\mathcal{X}\;\longrightarrow\;\mathbb{N}$
be defined by
$f^{z}(v)=n\quad\Longleftrightarrow\quad v\in\ell(n,z).$
Thus
$b(n,z)=\{v\;\Big{|}\;|v|_{M(R^{z}_{m})-n}\leq R^{z}_{m}\;for\;m\gg n\}$
Note that
$f^{z}(v+M(v))=f^{z}(v)+1$
and
$f^{z}(v+F(v))=f^{z}(v)+df^{z}(v),\qquad df^{z}(v)\in\{0,1\}.$
Thus $f^{z}$ is determined by $df^{z}$ taking values $0,1$, or by the subset
$Z(z)=\{\,v\in\mathcal{X}\;\mid\;df^{z}(v)=0\,\}.$
Moreover,
$Z(z)\cap\ell(n,z)=\bigl{\{}\,v\;\big{|}\;|v+F(v)|_{M(R^{z}_{m})-n}\;\leq\;R^{z}_{m}\;<\;|v|_{M(R^{z}_{m})-n+1},\;m\gg n\,\bigr{\}}.$
Normalization Lemma
Let $\mathbb{N}((t))$ denote the commutative rig (i.e. ring without negatives) of Laurent series with coefficients in $\mathbb{N}$.
Let
$\mathcal{B}:=\Bigl{\{}z(t)=\sum_{n}z_{n}t^{n}\in\mathbb{N}((t))\ \Big{|}\ V_{\varphi}(z):=\sum_{n}z_{n}\varphi^{n}<\infty\Bigr{\}}.$
This is a sub–rig of $\mathbb{N}((t))$, and we have a surjective rig homomorphism
$V_{\varphi}:\mathcal{B}\;\longrightarrow\;\mathbb{R}_{\geq 0}:=[0,\infty).$
We can view $z=(z_{n})\in\mathcal{F}$ as a power series $z(t)=\sum_{n}z_{n}t^{n},$
making $\mathcal{F}$ a subset of $\mathcal{B}$, and the $\varphi$–value is finite: $for\;z\in F,\;V_{\varphi}(z)=\sum_{n}z_{n}\varphi^{n}<\infty.$
We have a canonical projection
$P:\mathcal{B}\longrightarrow\mathcal{F},$
preserving the $\varphi$–value:
$V_{\varphi}(z)=V_{\varphi}(P(z)).$
For $z=(z_{n})\in\mathbb{N}^{\mathbb{Z}}$, with
$\sum_{n}z_{n}\varphi^{n}<\infty,$
(hence $z_{n}=0$ for $n\gg 0$), making the moves
  • (A)
    $(\dots 0\,z_{n}\,z_{n-1}\,z_{n-2}\dots)\;\;\longrightarrow\;\;(\dots 1\,(z_{n}-1)\,(z_{n-1}-1)\,z_{n-2}\dots)$
    if $z_{n}\geq 1$, $z_{n-1}\geq 1$,
  • (B)
[…]
    if $z_{n}\geq 2$,
produces a sequence that converges to $\mathcal{P}(z)\in\mathcal{F}$.
Note that steps (A), (B) preserve the $\varphi$-value.
Say that the sequence $z$ is OK above $z_{n}$ if
$z_{m}\in\{0,1\},\;\text{and}\;z_{m}\;\bullet\;z_{m-1}=0\;\text{ for all}\;m>n$
If we are in the case $z_{n}\geq 1$ and $z_{n-1}\geq 1$, a finite number of AA moves will produce a sequence that is OK above $z_{n-1}$, or a sequence OK above $z_{n}$, with $z_{n}\geq 2$, $z_{n-1}=0$. The last case is when move $B$ applies: it will replace $z_{n}$ by $z_{n}-2$, and after a few cleanup AA-moves it will be OK above $z_{n}$. Repeating this we see that given $(z)$ OK above $z_{n}$, it will be transformed to a sequence OK above $z_{n-1}$.
Note that for fixed $m\in\mathbb{Z}$, as $n\to$-$\infty$ the values of our sequence above $m$ can change, in which case the sum
$\sum_{j\geq m}z_{j}\varphi^{j}$
only increases, by an amount $\geq\varphi^{m}$.
The boundedness of the $\varphi$-value
$\sum_{n}z_{n}\varphi^{n}<\infty$
guarantees that this canonical process stabilizes above any finite index $m$.
Denoting by $P_{n}(z)$ the sequence obtained from $z$ after making it OK above $z_{n}$ in a fixed stable way, i.e. the $z_{m}$ for $m>n$ get their final value, so
$P_{n}(z)\xrightarrow[n\to-\infty]{}P(z)$
and for any finite index $m$, the coefficient
[…]
Rig structure on $\mathcal{F}$
Define for $z,z^{\prime}\in\mathcal{F}$:
$z+z^{\prime}:=P\!\bigl{(}z(t)+z^{\prime}(t)\bigr{)},\qquad z\cdot z^{\prime}:=P\!\bigl{(}z(t)\cdot z^{\prime}(t)\bigr{)}.$
These operations make $\mathcal{F}$ into a commutative rig, and $P$ is a rig homomorphism factoring $V_{\varphi}$:
$\mathcal{B}$$\mathbb{R}_{\geq 0}$$\mathcal{F}$$\scriptstyle{V_{\varphi}}$$\scriptstyle{P}$$\scriptstyle{V_{\varphi}}$
The operations $\bullet$ are clearly commutative, with unit for addition
$0=(z_{n}\equiv 0)\quad\text{the zero sequence,}$
and unit for multiplication
$1=(z_{n}\equiv\delta_{n})\quad\text{the sequence with }z_{0}=1,\;z_{n}=0\;\text{for }n\neq 0.$
To see the associativity of these operations, and the distributive law, one uses the identity
$P(x\circ y)=P(x\circ P(y)),\quad x,y\in B,$
with $\circ$ either addition or multiplication.
Notice that if $z^{\prime}$ is obtained from $z$ by a finite number of moves, each exchanging
$t^{n+1}\leftrightarrow t^{n}+t^{n-1},$
we have $P(z^{\prime})=P(z)$, and $P_{j}(z^{\prime})=P_{j}(z)$ for $j\ll 0$.
Thus
$P(x\circ y)=P(x\circ P_{n}(y)),\quad\text{all }n,$
but for fixed $m$, the final value of the $m$-coefficient involve only finitely many of the other coefficients, and so
$P(x\circ P_{n}(y))_{m}=P(x\circ P(y))_{m}$
for $n\ll N(m,V_{\varphi}(y),V_{\varphi}(x)).$
Let $\mathbb{N}[t,t^{-1}]\subseteq\mathcal{B}$ be the sub–rig generated by $t$ and $t^{-1}$.
Then $\mathcal{N}:=P(\mathbb{N}[t,t^{-1}])$ is the sub–rig of $\mathcal{F}$ consisting of finite sequences:
$\mathcal{N}=\Bigl{\{}z=(z_{n})\in\mathcal{F}\;\Big{|}\;z_{n}=0,\ z_{-n}=0\ \text{for }n\gg 0\Bigr{\}}.$
For $z\in\mathcal{N}\setminus\{0\}$, let $n_{0}=\delta(z)=\min\{n\;\mid\;z_{n}=1\}$, and define $z^{-}=(z_{n}^{-})\in\mathcal{F}$ by
[…]
Note that $z^{-}<z$, but they have the same $\varphi$–value:
$V_{\varphi}(z)=V_{\varphi}(z^{-}).$
The equivalence relation on $\mathcal{F}$ generated by
$z\sim z^{-}\qquad\text{for }z\in\mathcal{N}$
is compatible with the operations of addition and multiplication, and we have an identification
$\mathcal{F}\;\;\longrightarrow\;\;\mathcal{F}/\!\sim\;\;=\;\;\mathbb{R}_{\geq 0}$
Beware: Operations are not continuous on $\mathcal{F}$.
E.g.
$\lim_{n\to\infty}\bigl{(}1^{-}+\Phi^{-n}\bigr{)}=1\;>\;1^{-}=1^{-}+0=1^{-}+\lim_{n\to\infty}\Phi^{-n}.$
Quotients and Fundamental Domain
Let $\Phi\in\mathcal{F}$ be the image of $t$, so that $\Phi^{m}\cdot\Phi^{n}=\Phi^{m+n}$, and for $z\in\mathcal{F}$, $\Phi^{n}z$ is obtained by shifting its digits by $n$.
Note that
$R^{\Phi z}_{m}=R^{z}_{m+1}\;,b(n,\Phi z)=b(n,z)\;,f^{\Phi z}=f^{z}.$
Define:
$S:=\bigl{(}\mathcal{F}\setminus\{0\}\bigr{)}\big{/}\{\Phi^{n}\}_{n\in\mathbb{Z}}.$
Alternatively, let the ”Principal Units”
$\mathcal{D}:=\{\,z=(z_{n})\in\mathcal{F}\;\mid\;z_{0}=1\;\;\text{and}\;\;z_{n}=0\;\;\text{for }n>0\,\}.$
It is a fundamental domain for the action of $\{\Phi^{n}\}_{n\in\mathbb{Z}}$ on $\mathcal{F}\setminus\{0\}$, and we have an identification
$\mathcal{D}\;\;\longleftrightarrow\;\;S.$
We have multiplication on $S$, and taking $\varphi$–values gives a homomorphism onto the circle:
$V_{\varphi}:S\;\longrightarrow\;\mathbb{R}^{+}\big{/}\{\varphi^{n}\}_{n\in\mathbb{Z}}.$
We have a total order on $\mathcal{D}$, with minimal element $1=\Phi^{0}$, and maximal element
$(\Phi^{1})^{-}=(\ldots 0\ldots 01010101\ldots 01\ldots),$
and taking $\varphi$–values gives an order preserving map
$V_{\varphi}:\mathcal{D}\;\longrightarrow\;[1,\varphi].$
For $z\in\mathcal{D}$, we have
$\qquad M(R_{m}^{z})\equiv m$
and the base
$b(n,1)\;\subseteq\;b(n,z)\;\subseteq\;b\!\left(n,\Phi^{-}\right)$
Indeed for $z,z^{\prime}\in\mathcal{D}$
$z\leq z^{\prime}\;\;\text{implies}\;\;b(n,z)\subseteq b(n,z^{\prime}),\;and\;f^{z}(v)\geq f^{z^{\prime}}(v)\;\text{for all}\;v$
Let $B(n)$ denote the set of possible configurations of height $n$,
$B(n)=\{\,b(n,z)\mid z\in\mathcal{D}\,\}.$
We have restriction map $r:B(n)\to B(n-1)$, making
$\mathbb{B}=\bigcup_{n}B(n)$
into a planar rooted tree,
$\text{root: }B(1)=\{\,b(1,z)=\{(1,1)\}\;\;\text{any }z\in\mathcal{D}\,\}.$
Order on $B(n)$:
$b\leq b^{\prime}\;\;\Longleftrightarrow\;\;r^{j}b\;\subseteq\;r^{j}b^{\prime},\quad j=0,\ldots,n.$
The set $\mathcal{D}$ is the set of paths outgoing from the root of a planar binary rooted tree with vertices the finite $0,1$ expansions, with root $10$, and for a finite expansion $z$ we have its descendants $z0$ and $z10$.
The map
$\mathcal{D}\longrightarrow B(n),\quad z\longmapsto b(n,z)$
is continuous and order preserving.
It follows that there are unique $z^{n}_{j}\in\mathcal{N}=\mathbb{N}[\varphi,\varphi^{-1}]$
$1=z^{n}_{0}<z^{n}_{1}<\cdots<z^{n}_{k}<z^{n}_{k+1}<\cdots<z^{n}_{\#B(n)}=\Phi$
such that for $z\in\mathcal{D}$:
$z^{n}_{k}\leq z\leq(z^{n}_{k+1})^{-}\;\;\Longleftrightarrow\;\;b(n,z^{n}_{k})=b(n,z)=b\!\left(n,(z^{n}_{k+1})^{-}\right).$
Note that $\{z^{n-1}_{j}\}$ is a subset of $\{z^{n}_{k}\}$, and for $z^{n}_{k}\leq z\leq(z^{n}_{k+1})^{-}$
$r\,b(n,z)=b(n-1,z)=b\!\left(n-1,z^{n-1}_{j}\right)\quad\iff\quad[z^{n}_{k},(z^{n}_{k+1})^{-}]\subseteq[z^{n-1}_{j},(z^{n-1}_{j+1})^{-}].$
[…]
$f:X\to\mathbb{N},$
$f(v)=\min\{\,n\mid v\in b(n)\,\}$
The map
[…]
$z$ $\;\longmapsto\;\{b(n,z)\}_{n}\;\longleftrightarrow\;f^{z}$
is a bijection.
The Tree of configurations
We give the beginning of the tree $\mathbb{B}$ with the values of $z^{n}_{j}$’s, for n= 1,2,3,n\;=\;1,2,3, and with a denoting points of $Z(z)$ that determine the configuration, that is the points $v$ such that v+F(v)v\;+\;F(v) is in the same level as $v$, exhibiting the linear growth of $Z(z)$.
The root $b(1,z)$
123123(1,1)(1,1)
It branches to two in the first level
$b(2,z)$
$1=z^{2}_{0}\leq z<(z^{2}_{1})^{-}=10010\ldots 10\ldots$
$\text{i.e.}\;z=100\ldots$
$123$
$10100\ldots 0\ldots=z^{2}_{1}\leq z\leq z^{2}_{2}=\Phi^{-}$
$\text{i.e.}\;z=1010\ldots$
$123$
Each of these, branches to three in the next level.
$b(3,z)$
$z=100\ldots$
$123$
$1=z^{3}_{0}\leq z$
$z\leq(z^{3}_{1})^{-}=10000010\ldots 10\ldots$
$1234$
$1000010\ldots 0\ldots\;=\;z^{3}_{1}\;\leq\;z$
$z\leq\;(z^{3}_{2})^{-}\;=\;10010010\ldots 10\ldots$
$12345$
$10010100\ldots 0\ldots\;=\;z^{3}_{2}\;\leq\;z$
$z\;\leq\;(z^{3}_{3})^{-}=\;10010\ldots 10\ldots$
$z=101\ldots$
$12345$
$\text{1010}\ldots 0\ldots\;=\;z^{3}_{3}\;\leq\;z$
$z\leq\;(z^{3}_{4})^{-}\;=\;10100010\ldots 10\ldots$
$12345$
$1010010\ldots 0\ldots\;=\;z^{3}_{4}\;\leq\;z$
$z\leq\;(z^{3}_{5})^{-}\;=\;10101000\,10\ldots 10\ldots$
$123456$
$10101001000\ldots 0\ldots\;=\;z^{3}_{5}\;\leq\;z$
$z\leq\;(z^{3}_{6})^{-}\;=\;\Phi^{-}\;=\;1010\ldots 10\ldots$